# Advancing Grounded Multimodal Named Entity Recognition via LLM-Based Reformulation and Box-Based Segmentation

### v1

@@ Abstract
Grounded Multimodal Named Entity Recognition (GMNER) task aims to identify named entities, entity types and their corresponding visual regions. GMNER task exhibits two challenging attributes: 1) The tenuous correlation between images and text on social media contributes to a notable proportion of named entities being ungroundable. 2) There exists a distinction between coarse-grained noun phrases used in similar tasks (e.g., phrase localization) and fine-grained named entities. In this paper, we propose RiVEG, a unified framework that reformulates GMNER into a joint MNER-VE-VG task by leveraging large language models (LLMs) as connecting bridges. This reformulation brings two benefits: 1) It enables us to optimize the MNER module for optimal MNER performance and eliminates the need to pre-extract region features using object detection methods, thus naturally addressing the two major limitations of existing GMNER methods. 2) The introduction of Entity Expansion Expression module and Visual Entailment (VE) module unifies Visual Grounding (VG) and Entity Grounding (EG). This endows the proposed framework with unlimited data and model scalability. Furthermore, to address the potential ambiguity stemming from the coarse-grained bounding box output in GMNER, we further construct the new Segmented Multimodal Named Entity Recognition (SMNER) task and corresponding Twitter-SMNER dataset aimed at generating fine-grained segmentation masks, and experimentally demonstrate the feasibility and effectiveness of using box prompt-based Segment Anything Model (SAM) to empower any GMNER model with the ability to accomplish the SMNER task. Extensive experiments demonstrate that RiVEG significantly outperforms SoTA methods on four datasets across the MNER, GMNER, and SMNER tasks. Datasets and Code will be released at https://github.com/JinYuanLi0012/RiVEG.
@@
Multimodal Named Entity Recognition (MNER) on social media is a classical multimodal information extraction task [1, 2]. Due to the abundance of informal and brief unstructured textual content on social media platforms, many studies [3, 4, 5] attempt to improve Named Entity Recognition (NER) performance by leveraging multimodal features. Compared with traditional NER methods that solely consider text information [6, 7, 8, 9, 10], MNER approaches demonstrate superior performance in many scenarios [11, 12, 13, 14, 15, 16]. However, MNER only focuses on extracting entity-type pairs from image-text pairs and cannot fully address the pressing needs in related domains, such as building multimodal knowledge graphs and enhancing human-computer interaction. Recent studies [17] endeavor to conduct more in-depth extensions on MNER. As an emerging multimodal task, Grounded Multimodal Named Entity Recognition (GMNER) aims to extract text named entities and entity types from image-text pairs while performing visual grounding of named entities in form of bounding boxes.
@@
The entire framework is endowed with unlimited data scalability and model scalability by reformulating the task. On the one hand, the existing 7k training data is limited for the realistic and challenging task like GMNER. However, once the GMNER task is reframed as a joint composition of MNER-VE-VG modules, the entire framework can naturally inherit the corresponding pre-training foundations of different modules, e.g., about 6.1M language expressions and 174k images [23, 29, 30, 31, 32] available for the VG module pre-training [24], about 14.1M image-text pairs [33, 34, 35, 36, 37] available for the VE module pre-training [28], etc. Furthermore, since different LLMs provide diverse auxiliary knowledge and entity expansion expressions for the same sample, the entire framework can further use various LLMs to enhance the limited 7k training data to gain data marginal benefits. On the other hand, wider training data supports better model scalability. The entire framework can naturally adapt to the relevant research and development of any single module to gain more model architecture advantages, and can freely select more accurate or lightweight independent modules based on application scenarios to constitute various variants. Experimental results show that all 14 variants of RiVEG exhibit better performance compared with existing baseline methods.
@@
The new task and dataset are constructed to achieve more fine-grained multimodal information extraction with almost no performance loss. Since generating fine-grained segmentation masks of visual objects can effectively solve the ambiguity that may result from output bounding boxes, we propose a new Segmented Multimodal Named Entity Recognition (SMNER) task and construct a dataset that can be used for SMNER task by manually annotating the visual object segmentation masks of all groundable named entities in existing GMNER dataset. The SMNER task is more challenging than the GMNER task, because the former not only requires predicting the groundability of each entity but also predicting the fine-grained segmentation mask, which is more difficult than predicting the coarse-grained bounding box. Although it is challenging to directly build a SMNER model with acceptable performance based on about 4k groundable named entity samples in this dataset, the emergence of Segment Anything Model (SAM) [38] makes it possible to solve this task. SAM cannot identify masked objects based on arbitrary text input. But considering that benefiting from prior knowledge of box positions and predicting the corresponding object masks based on bounding boxes can be more effective than directly predicting region masks based on text [39], we integrate the capabilities of GMNER methods and SAM. With the entity groundability judgment and visual localization capabilities already possessed by GMNER methods, as well as the outstanding zero-shot segmentation ability of SAM based on box prompts, we demonstrate that any GMNER method can be equipped with the capability to accomplish SMNER task with almost no performance penalty.111This paper is a substantial extension of our previous Findings of EMNLP 2023 [40] and ACL 2024 [41] works. There are three major improvements in this journal version, which are detailed in the Supplementary Material. And these will be added to the end of the introduction in the revised version.
@@ II Background
II-A Taxonomy of MNER Methods
@@
II-B Existing GMNER Methods
@@
The existing GMNER method uses object detection (OD) methods [46, 47] to detect all visual objects in images, and then performs entity-region matching between named entities and all candidate visual objects. But as shown in Table I, these classic OD methods perform poorly due to the complexity and diversity of images originating from social media.222Here, TopN-Prec@0.5 is the proportion of entities where at least one of the Top-N predicted bounding boxes based on detection probability has an IoU of 0.5 or greater with the ground truth bounding box. This means that there is a natural performance ceiling for this kind of matching. On the one hand, samples that fail to detect ground truth visual regions by OD methods in the prediction stage inevitably have erroneous entity-region matching. On the other hand, the existing method treat named entity samples where ground truth visual regions are not successfully detected by OD methods as ungroundable samples in the training stage (i.e., 7088/4694 to 8262/3520 ungroundable/groundable). This exacerbates the data imbalance of the GMNER dataset [17], potentially leading the model to bias towards predicting entities as ungroundable. Therefore, it is necessary to avoid such entity-region matching based on OD methods.
@@ III-A Dataset Construction
Since no dataset exists for the SMNER task, we construct a Twitter-SMNER dataset as follows. The existing GMNER dataset [17] accomplishes data collection, cleaning, and coarse-grained bounding box annotation of named entities. We further annotate fine-grained segmentation masks based on this foundation. Specifically, we employ five annotators and two experienced experts, and standardize the annotation tool and rules.333https://github.com/labelmeai/labelme Each image-text pair is assigned to two different annotators. For controversial annotation samples, decisions are made by experienced experts. A tiny fraction of ambiguous samples is removed during the annotation process. Two annotations with IoU greater than 0.5 are considered consistent annotations. The Fleiss score [48] between two different annotators is 0.82, indicating a high level of consistency.
@@ III-B Dataset Analysis
We maintain the same partitioning as the Twitter-GMNER dataset [17], i.e., 70% is used for training set, 15% for dev set, and 15% for test set. As shown in Table II, the Twitter-SMNER dataset contains 16778 text named entities. For the 6681 groundable named entities, we annotate 7973 segmentation masks. This means that one entity may correspond to multiple segmentation masks. Fig. 4 (left) shows that 43.9% of the images do not contain any segmentation mask, 42.4% of the images contain a single segmentation mask, and about 13.7% of the images contain multiple segmentation masks. And Fig. 4 (right) shows that named entities of the PER type are often present in images, while the other three types are not. This reflects the challenge of the proposed SMNER task.
@@ IV Methodology
RiVEG mainly consists of two stages. The first stage aims to ensure optimal MNER performance and further perform entity expansion expression (detailed in § IV-B). The second stage aims to reformulate the entire EG task as a joint of VE and VG (detailed in § IV-C). We also conduct direct and effective data augmentation on the overall framework using different LLMs. The overall architecture of RiVEG is shown in Fig. 5.
@@ IV-A Task Formulation
where represents the text named entity, represents the entity type of (i.e., PER, LOC, ORG and MISC), represents the visually grounded region of , and represents the visual segmentation mask of . Note that if has no corresponding visual object in the image, or is specified as None. Otherwise, comprises 4D coordinates representing the top-left and bottom-right positions of the bounding box, while consists of the binary segmentation mask. For predicted visual region and segmentation mask , it is considered correct only if its IoU with a certain gold label is greater than 0.5. For all three tasks, only predictions where all elements are correct are considered correct.
@@ IV-B2 Entity Expansion Expression Module
In order to bridge the difference between real-world named entities and noun phrases, we design the Entity Expansion Expression module to further guide LLMs to reformulate fine-grained named entities into coarse-grained entity expansion expressions. Specifically, we formulate all named entities into the following template:
@@
where is the answer generated by LLMs. Different from the process of generating auxiliary refined knowledge, since the length of entity expansion expression is much shorter than that of auxiliary refined knowledge, the awareness of similar examples is no longer involved here. Instead, a small set of fixed in-context examples is employed. Finally, we concatenate named entities, entity types, and entity expansion expressions generated by LLMs in the following format to form the final named entity referring expressions: . Likewise, different LLMs are utilized to generate diverse expansion expressions for the same named entity, effectively achieving data augmentation.
@@
Through the above processing, the entire GMNER task can be naturally redefined as the MNER-VE-VG task. Due to the abundance of existing research on VE [51, 28, 27, 52] and VG [24, 25, 26, 27], the method at this stage is also not unique. Note that this study emphasizes the coordination between various modules rather than alterations in the model architecture. Because one of our primary motivations is to explore how to effectively inherit the pre-training foundation of related research in the GMNER task with very limited training data, thereby giving the entire framework unlimited model scalability and data scalability.
@@ IV-C1 Visual Entailment Module
Our experimental results show that when SoTA VG methods are directly fine-tuned using the Twitter-GMNER training set [17], these methods achieve Prec@0.5 scores of 21.87% [27] and 9.23% [24] on the Twitter-GMNER test set.444For the fine-tuning process, we specify the text input as the named entity along with the entity type. For the ungroundable text input, we define the 4D gold label as (). This demonstrates the difficulty of endowing existing VG methods with the capability to handle the ungroundable text input through direct fine-tuning with limited GMNER data. To address the disparity between VG methods and the EG task, we introduce the Visual Entailment (VE) module. Classic VE task [51] defines the VE dataset as:
@@
where is the image, is the named entity referring expression, and is the class label. Since the groundability of each text named entity in the GMNER dataset is explicit, retains only two labels and .555Here, and are not distinguished and can be considered the same. This is because the Twitter-SMNER dataset is constructed based on the Twitter-GMNER dataset. There is no significant difference in the groundability of named entities in these two datasets. represents that can be grounded in , and represents that cannot be grounded in . During the training phase, the dataset defined above is utilized to fine-tune existing VE models. And during the inference phase, only named entity referring expressions that are considered groundable by the fine-tuned VE model are input to the subsequent VG module. Ungroundable named entity referring expressions are filtered and specified with their entity grounding result as None.
@@ IV-C2 Visual Grounding and Segmentation Module
Additionally, because endowing the model with the ability to output precise pixel-level object masks is more challenging than endowing the model with the ability to output visual object 4D coordinates, and the Twitter-SMNER training set contains less than 5k groundable named entities, it is difficult to obtain satisfactory performance by migrating the above processing method for the VG models directly to the Reference Expression Segmentation models [26, 24, 53].666Table X shows the detailed experimental test results. However, a series of previous works [54, 55] show that it is feasible to use the box position prompt to obtain the segmentation mask of the main visual object in the box, and previous work such as OpenSeeD [39] also shows that obtaining the region mask based on the box position prompt is more effective than predicting the region mask directly based on the text query. We instead try to combine the localization capability of the VG method and the zero-shot image segmentation ability of the SAM [38]. Specifically, for any input sample in the Twitter-SMNER dataset, the VE method trained on the is used to determine the groundability of the named entity referring expression . For determined to be groundable, the VG method trained on is used to predict the 4D coordinates of the visual object associated with , i.e., (). SAM receives the 4D coordinates and treats that coordinates as the box prompt, generating the segmentation mask of :

### v2

@@ Abstract
Grounded Multimodal Named Entity Recognition (GMNER) task aims to identify named entities, entity types and their corresponding visual regions. GMNER task exhibits two challenging attributes: 1) The tenuous correlation between images and text on social media contributes to a notable proportion of named entities being ungroundable. 2) There exists a distinction between coarse-grained noun phrases used in similar tasks (e.g., phrase localization) and fine-grained named entities. In this paper, we propose RiVEG, a unified framework that reformulates GMNER into a joint modeling paradigm spanning MNER, VE, and VG perspectives by leveraging large language models (LLMs) as connecting bridges. This reformulation brings two benefits: 1) It enables us to optimize the MNER module for optimal MNER performance and eliminates the need to pre-extract region features using object detection methods, thus naturally addressing the two major limitations of existing GMNER methods. 2) The introduction of Entity Expansion Expression module and Visual Entailment (VE) module unifies Visual Grounding (VG) and Entity Grounding (EG). This endows the proposed framework with unlimited data and model scalability. Furthermore, to address the potential ambiguity stemming from the coarse-grained bounding box output in GMNER, we further construct the new Segmented Multimodal Named Entity Recognition (SMNER) task and corresponding Twitter-SMNER dataset aimed at generating fine-grained segmentation masks, and experimentally demonstrate the feasibility and effectiveness of using box prompt-based Segment Anything Model (SAM) to empower any GMNER model with the ability to accomplish the SMNER task. Extensive experiments demonstrate that RiVEG significantly outperforms SoTA methods on four datasets across the MNER, GMNER, and SMNER tasks. Datasets and Code will be released at https://github.com/JinYuanLi0012/RiVEG.
@@
Multimodal Named Entity Recognition (MNER) on social media is a classical multimodal information extraction task [1, 2]. Due to the abundance of informal and brief unstructured textual content on social media platforms [3, 4, 5], many studies [6, 7, 8] attempt to improve Named Entity Recognition (NER) performance by leveraging multimodal features. Compared with traditional NER methods that solely consider text information [9, 10, 11, 12], MNER approaches demonstrate superior performance in many scenarios [13, 14, 15, 16, 17, 18]. However, MNER only focuses on extracting entity-type pairs from image-text pairs and cannot fully address the pressing needs in related domains, such as building multimodal knowledge graphs [19] and enhancing human-computer interaction [20]. Recent studies [21, 22] endeavor to conduct more in-depth extensions on MNER. As an emerging multimodal task, Grounded Multimodal Named Entity Recognition (GMNER) aims to extract text named entities and entity types from image-text pairs while performing visual grounding of named entities in form of bounding boxes.
@@
The entire framework is endowed with unlimited data scalability and model scalability by reformulating the task. On the one hand, the existing 7k training data is limited for the realistic and challenging task like GMNER. However, once the GMNER task is reframed as a joint composition of modules grounded in MNER, VE, and VG, the entire framework can naturally inherit the corresponding pre-training foundations of different modules, e.g., about 6.1M language expressions and 174k images [28, 34, 35, 36] available for the VG module pre-training [29], about 14.1M image-text pairs [37, 38, 39, 40, 41] available for the VE module pre-training [33], etc. Furthermore, since different LLMs provide diverse auxiliary knowledge and entity expansion expressions for the same sample, the entire framework can further use various LLMs to enhance the limited 7k training data to gain data marginal benefits. On the other hand, wider training data supports better model scalability. The entire framework can naturally adapt to the relevant research and development of any single module to gain more model architecture advantages, and can freely select more accurate or lightweight independent modules based on application scenarios to constitute various variants. Experimental results show that all 14 variants of RiVEG exhibit better performance compared with existing baseline methods.
@@
The new task and dataset are constructed to achieve more fine-grained multimodal information extraction with almost no performance loss. Since generating fine-grained segmentation masks of visual objects can effectively solve the ambiguity that may result from output bounding boxes, we propose a new Segmented Multimodal Named Entity Recognition (SMNER) task and construct a dataset that can be used for SMNER task by manually annotating the visual object segmentation masks of all groundable named entities in existing GMNER dataset. The SMNER task is more challenging than the GMNER task, because the former not only requires predicting the groundability of each entity but also predicting the fine-grained segmentation mask, which is more difficult than predicting the coarse-grained bounding box. Although it is challenging to directly build a SMNER model with acceptable performance based on about 4k groundable named entity samples in this dataset, the emergence of Segment Anything Model (SAM) [42] makes it possible to solve this task. SAM cannot identify masked objects based on arbitrary text input. But considering that benefiting from prior knowledge of box positions and predicting the corresponding object masks based on bounding boxes can be more effective than directly predicting region masks based on text [43], we integrate the capabilities of GMNER methods and SAM. With the entity groundability judgment and visual localization capabilities already possessed by GMNER methods, as well as the outstanding zero-shot segmentation ability of SAM based on box prompts, we demonstrate that any GMNER method can be equipped with the capability to accomplish SMNER task with almost no performance penalty.
This paper is a substantial extension of our previous Findings of EMNLP 2023 [44] and ACL 2024 [45] works. This journal version has three major improvements:
We propose a new Segmented Multimodal Named Entity Recognition (SMNER) task, construct the corresponding Twitter-SMNER dataset, and provide detailed information about its construction process.
To complete and evaluate the SMNER task, we effectively extend three existing GMNER methods [21] and the proposed RiVEG framework [45]. By leveraging the box prompt-based SAM [42], we demonstrate the feasibility of enabling any existing GMNER method to complete the SMNER task. This also allows us to discuss the adaptability of two different segmentation mask acquisition methods (i.e., text query-based and box prompt-based) for the SMNER task.
We present more detailed and comprehensive quantitative and qualitative experiments on the PLIM module [44] and the RiVEG framework [45], including evaluations with various LLMs and text encoders, IoU threshold sensitivity analyses for the GMNER and SMNER tasks, as well as clearer and more intuitive case visualizations.
@@ II Background
II-A Terminology and Abbreviations
Given the complexity of the proposed framework and the variety of tasks and modules involved, we summarize the main abbreviations in Table I for clarity. These terms will appear frequently in the remainder of the paper.
| Abbreviation | Full Term |
| NER | Named Entity Recognition |
| MNER | Multimodal Named Entity Recognition |
| GMNER | Grounded Multimodal Named Entity Recognition |
| SMNER | Segmented Multimodal Named Entity Recognition |
| LLM | Large Language Model |
| OD | Object Detection |
| VG | Visual Grounding |
| VE | Visual Entailment |
| SAM | Segment Anything Model |
| PLIM | Prompt Large Language Models In MNER |
| MSEA | Multimodal Similar Example Awareness |
| RiVEG | Reformulating GMNER into MNER, VE, and VG |
II-B Taxonomy of MNER Methods
@@
II-C Existing GMNER Methods
@@
The existing GMNER method uses object detection (OD) methods [50, 51] to detect all visual objects in images, and then performs entity-region matching between named entities and all candidate visual objects. But as shown in Table II, these classic OD methods perform poorly due to the complexity and diversity of images originating from social media.111Here, TopN-Prec@0.5 is the proportion of entities where at least one of the Top-N predicted bounding boxes based on detection probability has an IoU of 0.5 or greater with the ground truth bounding box. This means that there is a natural performance ceiling for this kind of matching. On the one hand, samples that fail to detect ground truth visual regions by OD methods in the prediction stage inevitably have erroneous entity-region matching. On the other hand, the existing method treat named entity samples where ground truth visual regions are not successfully detected by OD methods as ungroundable samples in the training stage (i.e., 7088/4694 to 8262/3520 ungroundable/groundable). This exacerbates the data imbalance of the GMNER dataset [21], potentially leading the model to bias towards predicting entities as ungroundable. Therefore, it is necessary to avoid such entity-region matching based on OD methods.
@@ III-A Dataset Construction
Since no dataset exists for the SMNER task, we construct a Twitter-SMNER dataset as follows. The existing GMNER dataset [21] accomplishes data collection, cleaning, and coarse-grained bounding box annotation of named entities. We further annotate fine-grained segmentation masks based on this foundation. Specifically, we employ five annotators and two experienced experts with research backgrounds in computer vision. The annotation tool and guidelines are also standardized to ensure consistency.222https://github.com/labelmeai/labelme Each image-text pair is assigned to two different annotators. For controversial annotation samples, decisions are made by experienced experts. A tiny fraction of ambiguous samples is removed during the annotation process. Two annotations with IoU greater than 0.5 are considered consistent annotations. The Fleiss score between two different annotators is 0.82 [52]. To further assess the consistency of the annotations, we also calculate the Dice Coefficient [53], which yields a score of 0.85. These results collectively demonstrate the high annotation consistency of the Twitter-SMNER dataset.
In the data annotation process, we follow the standard visible instance segmentation setting that is widely adopted in datasets such as COCO [39] and LVIS [54]. As shown in Fig. 4, each groundable entity is annotated with a separate segmentation mask that covers only the visible (non-occluded) regions. When occlusion occurs between different entities, their segmentation masks are allowed to overlap.
@@ III-B Dataset Analysis
We maintain the same partitioning as the Twitter-GMNER dataset [21], i.e., 70% is used for training set, 15% for dev set, and 15% for test set. As shown in Table III, the Twitter-SMNER dataset contains 16778 text named entities. For the 6681 groundable named entities, we annotate 7973 segmentation masks. This means that one entity may correspond to multiple segmentation masks. Fig. 5 (left) shows that 43.9% of the images do not contain any segmentation mask, 42.4% of the images contain a single segmentation mask, and about 13.7% of the images contain multiple segmentation masks. And Fig. 5 (right) shows that named entities of the PER type are often present in images, while the other three types are not. This reflects the challenge of the proposed SMNER task.
@@ IV Methodology
RiVEG mainly consists of two stages, aligned with the pipeline design established in our previous conference paper [44, 45]. The first stage aims to ensure optimal MNER performance and further perform entity expansion expression (detailed in § IV-B). The second stage aims to reformulate the entire EG task as a joint of VE and VG (detailed in § IV-C). Furthermore, this journal version extends RiVEG by incorporating a new segmentation head in the second stage to support the SMNER task, where bounding-box prompts are used to guide the generation of segmentation masks. In addition, we employ various LLMs to perform direct and effective data augmentation on the overall framework in this extended version. The overall architecture of RiVEG is shown in Fig. 6.
@@ IV-A Task Formulation
where represents the text named entity, represents the entity type of (i.e., PER, LOC, ORG and MISC), represents the visually grounded region of , and represents the visual segmentation mask of . Note that if has no corresponding visual object in the image, or is specified as None. Otherwise, comprises 4D coordinates representing the top-left and bottom-right positions of the bounding box, while consists of the binary segmentation mask. For predicted visual region and segmentation mask , it is considered correct only if its IoU with a certain gold label is greater than 0.5. For the SMNER task, the pixel-level IoU metric is adopted, which directly reflects the overlap between the predicted mask and the ground truth mask. For the GMNER task, the box-level IoU metric is employed, which measures the overlap between the predicted bounding box and the ground truth box. For all three tasks, only predictions where all elements are correct are considered correct.
@@ IV-B2 Entity Expansion Expression Module
To bridge the gap between real-world named entities and noun phrases, we design an Entity Expansion Expression module to guide LLMs in reformulating fine-grained named entities into semantically meaningful coarse-grained expressions. Specifically, we structure the input to LLMs using the following prompt template:
@@
Here, the is generated by LLMs. Unlike the generation of auxiliary refined knowledge, which requires dynamic awareness of similar examples due to its length and complexity, the generation of expansion expressions relies on a fixed set of manually constructed in-context examples. This is because the target expansion expressions are short and concise, making dynamic selection less impactful and more computationally expensive.
Finally, we concatenate the named entity, its entity type, and the generated entity expansion expression to form the final named entity referring expression in the format: . To promote diversity and enable data augmentation, we utilize different LLMs to produce multiple expansion expressions for the same named entity.
@@
Through the above processing, the entire GMNER task can be naturally redefined as a joint modeling paradigm spanning the perspectives of MNER, VE, and VG. Due to the abundance of existing research on VE [56, 33, 32] and VG [29, 30, 31, 32], the method at this stage is also not unique. Note that this study emphasizes the coordination between various modules rather than alterations in the model architecture. Because one of our primary motivations is to explore how to effectively inherit the pre-training foundation of related research in the GMNER task with very limited training data, thereby giving the entire framework unlimited model scalability and data scalability.
@@ IV-C1 Visual Entailment Module
Our experimental results show that when SoTA VG methods are directly fine-tuned using the Twitter-GMNER training set [21], these methods achieve Prec@0.5 scores of 21.87% [32] and 9.23% [29] on the Twitter-GMNER test set.333For the fine-tuning process, we specify the text input as the named entity along with the entity type. For the ungroundable text input, we define the 4D gold label as (). This demonstrates the difficulty of endowing existing VG methods with the capability to handle the ungroundable text input through direct fine-tuning with limited GMNER data. To address the disparity between VG methods and the EG task, we introduce the Visual Entailment (VE) module. Classic VE task [56] defines the VE dataset as:
@@
where is the image, is the named entity referring expression, and is the class label. Since the groundability of each text named entity in the GMNER dataset is explicit, retains only two labels and .444Here, and are not distinguished and can be considered the same. This is because the Twitter-SMNER dataset is constructed based on the Twitter-GMNER dataset. There is no significant difference in the groundability of named entities in these two datasets. represents that can be grounded in , and represents that cannot be grounded in . During the training phase, the dataset defined above is utilized to fine-tune existing VE models. And during the inference phase, only named entity referring expressions that are considered groundable by the fine-tuned VE model are input to the subsequent VG module. Ungroundable named entity referring expressions are filtered and specified with their entity grounding result as None.
@@ IV-C2 Visual Grounding and Segmentation Module
Additionally, because endowing the model with the ability to output precise pixel-level object masks is more challenging than endowing the model with the ability to output visual object 4D coordinates, and the Twitter-SMNER training set contains less than 5k groundable named entities, it is difficult to obtain satisfactory performance by migrating the above processing method for the VG models directly to the Reference Expression Segmentation models [31, 29].555Table XI shows the detailed experimental test results. However, a series of previous works [57, 58] show that it is feasible to use the box position prompt to obtain the segmentation mask of the main visual object in the box, and previous work such as OpenSeeD [43] also shows that obtaining the region mask based on the box position prompt is more effective than predicting the region mask directly based on the text query. We instead try to combine the localization capability of the VG method and the zero-shot image segmentation ability of the SAM [42]. Specifically, for any input sample in the Twitter-SMNER dataset, the VE method trained on the is used to determine the groundability of the named entity referring expression . For determined to be groundable, the VG method trained on is used to predict the 4D coordinates of the visual object associated with , i.e., (). SAM receives the 4D coordinates and treats that coordinates as the box prompt, generating the segmentation mask of :